# Loss of Control Prevention of an Agile Aircraft: Dynamic Command Saturation Approach

### v1

@@
Ege Ç. Altunkaya  [^1] Research Assistant, Aviation Institute, Aerospace Research Center, Istanbul Technical University, Istanbul, Türkiye, 34469. and Akın Çatak  [^2] Research Assistant, Aviation Institute, Aerospace Research Center, Istanbul Technical University, Istanbul, Türkiye, 34469. Aerospace Research Center, Istanbul Technical University, Istanbul, Türkiye, 34469 Emre Koyuncu  [^3] Assoc. Prof., Aerospace Research Center, Istanbul Technical University, Istanbul, Türkiye, 34469. Aerospace Research Center, Istanbul Technical University, Istanbul, Türkiye, 34469 İbrahim Özkol  [^4] Prof., Aerospace Research Center, Istanbul Technical University, Istanbul, Türkiye, 34469. Aerospace Research Center, Istanbul Technical University, Istanbul, Türkiye, 34469
@@ 1 Introduction
The loss of control, a leading factor in fatal crashes, poses a critical threat to flight safety by jeopardizing the flight stability [1]. Numerous factors contribute to this loss, ranging from adverse onboard conditions to unexpected vehicle upsets [2, 3]. The diverse array of reasons and their interactions inherently elevate the risk of control loss, as extensively outlined in terms of fatal crash statistics linked to the control loss in [2] and [3]. Moreover, a detailed analysis of the causes behind the control loss reveals that the pilot-induced loss of control, a significant contributing factor, involves abrupt and aggressive maneuvers that can trigger a destabilized flight [3]. Despite the desirability of performing aggressive maneuvers without compromising safety, it remains constrained by the aircraft’s aerodynamic characteristics, controllability, and its flight dynamics. Additionally, according to [4], reasons for loss of control can be categorized under five major headlines: stall, sideslip-induced rolls, rolls from other causes, pilot-induced oscillation, and yaw.
Thereby designing an aircraft with a high-performance flight control law is not always sufficient and does not imply that the aircraft can perform any sort of maneuver. There are numerous restrictions that hinder an agile aircraft from conducting agile maneuvers, such as high angle of attack kinematics, inertia-coupling, and flight and/or dynamic envelope violation. The preceding factors are so hazardous that they may trigger an upset condition such as loss of control, departure, spin, stall, or even structural destruction. In summary, loss of control refers to a flight dynamics and control issue where the aircraft’s motion deviates from expected operating conditions, and this problem involves non-linear effects like nonlinear kinematic and inertial coupling, which cause enormous responses to small changes in state variables or result in the oscillatory or divergent behavior. As a consequence, it often leads to significant angular rates and displacements, making it challenging to maintain the desired heading, altitude, and level flight [5]. Therefore, to prevent any upset condition, one has to delve into presumable reasons for these catastrophic consequences.
@@
Since loss of control still persists as a critical issue, numerous studies inherently exist to address it from various perspectives. Hereby, ensuring safe flight has been extensively studied through flight envelope protection [6, 7, 8, 9, 10, 11, 12, 13, 14], prevention of loss of control [15, 9, 16], and implementation of command/reference governor methods [17, 18]. Lombaerts et al. [9] detail the design process of the adaptive algorithm and its integration into a piloted simulator for evaluation. The paper discusses the methodology employed in simulator testing, including the scenarios used to assess the algorithm’s performance. The boundaries of the flight envelope are determined adaptively, and the flight envelope-aware flight control system serves to prevent loss of control. Finally, they implement the proposed architecture into a simulator environment, belonging to a commercial airline aircraft, for further examination. Yu et al. [13] present an approach for designing safety control systems for aircraft that incorporate flight envelope protection and reference command generation. The approach is based on a nonlinear control framework that combines feedback control with feedforward reference commands to achieve robust and safe flight performance under a variety of operating conditions. The results show that the approach is effective at maintaining stable flight and protecting the aircraft from operating outside of its safe flight envelope. Despite the existence of adaptive flight envelope determination methods, there also exist other online safe flight envelope estimation studies [19, 20] besides the ANN-based flight envelope estimation approach [21].
@@
Nevertheless, various studies strongly emphasize the distinction between the flight envelope and dynamic envelope of an aircraft [22, 1]. In other words, the dynamic envelope defines the region within the aircraft’s state space where safe controllability is assured, preventing the occurrence of events leading to loss of control. Consequently, the notion of ensuring flight conditions solely within the traditional flight envelope for the prevention of the loss of control becomes a matter of debate. Particularly, Zheng et al. [22] study the determination of the dynamic envelope of the baseline aircraft, NASA’s Generic Transport Model. They generate the dynamic envelope using differential manifold theory and claim that this method is able to capture the coupling effects more accurately. Moreover, a correlation between the dynamic envelope and control coefficient or pilot command is established. Finally, based on the dynamic envelope enlargement strategy, an envelope protection architecture is proposed and assessed.
@@
On the other hand, various studies aim to estimate the maneuvering envelope through reachability analysis, such as [23, 24, 15, 20]. However, these approaches are not suitable for the real-time adaptation due to the high computational cost associated with numerically solving the Hamilton-Jacobi-Bellman equations. Furthermore, in [1, 25], loss of control is thoroughly investigated in the sense of nonlinear dynamics, and bifurcation analysis is put forth. However, of course, bifurcation analyses can only be done in the design phase of the aircraft.
@@
Last but not least, maneuver limiting is also addressed in various studies using the command and reference governor approaches. Simon et al. [17] propose a method for limiting the angle of attack and load factor using the command and reference governors. The implementation has been carried out in the simulation environment used for the Jas39 Gripen fighter aircraft. As a result of the study, a remarkable degradation in overshoot is observed in the angle of attack and load factor responses at their limit values, constrained by the corresponding flight envelope. A similar effort has been undertaken by [18] for the flight envelope protection during high-angle-of-attack maneuvers. Albeit the proposed methods are successful in achieving their primary objectives, i.e., flight envelope protection, a priori flight envelope information is a must; otherwise, command and reference governors cannot determine where to stop the aircraft prior to entering an upset condition.
@@
From an industrial perspective, the primary goal is "care-free" maneuvering without compromising performance and safety, as emphasized in [25]. However, the certification remains a constant requirement in the industry, making conventional methods the preferred choice for safety. Traditional state limiters, such as angle of attack, load factor, bank angle, pitch angle, and roll rate limiters, are still in use. These limiters are scheduled based on the flight-critical parameters like Mach number, dynamic pressure, angle of attack, and/or roll rate. Nevertheless, designing these state limiters for safe flight requires a labor-intensive effort. Extensive aerodynamic analyses are necessary to investigate the stability characteristics of the designed aircraft. For this purpose, several helpful notions (not limited to these), such as lateral control departure parameter and $C_{{n_{\beta}}_{dyn}}$, are adopted. Careful analysis of these parameters is of significant importance to determine the departure angle of attack value, and their variations with changing flight circumstances must be considered for scheduling. Importantly, these evaluations are not limited to the early design phases but must be extended and verified through the flight tests. Consequently, these aspects contribute to the labor and cost excessively.
@@
According to the inferences drawn from the previous research, as elaborated previously, the issue of the loss of control has been discussed within the context of the flight envelope protection. Additionally, a priori flight envelope databases generally exist in these studies. Consequently, in such investigations, the aircraft is pre-aware of where to avoid based on the pre-known flight envelope. However, confining the concept of the loss of control solely within the flight envelope protection does not seem inclusive, and the consideration of extreme and coupling maneuvers has been lacking in most previous research. Based on these evaluations, the followings are the contributions of this study:
  • A priori information about where to avoid from does not exist in this study; however, this information is derived in an online manner by merging the Lyapunov’s direct method and inverse dynamics of the aircraft.
  • Controllability investigation has never been handled using the incremental attainable moment set concept. Both the derivation of the incremental attainable moment set and its utilization for an online controllability assessment are discussed.
  • Considerably extreme maneuvers are evaluated to observe the limits of the aircraft by incorporating the coupling and nonlinear effects, contrary to what has been done in the previous works. The efficacy of the proposed method is demonstrated based on these harsh maneuvers.
@@
  • Significantly more ample stable maneuverable volumes are acquired as a result of the comparison between the conventional state limiters and the proposed approach, which is also a pure indicator of the enhancement of the aircraft’s maneuverability without compromising safety.
@@
Consequently, the paramount importance of this study lies in its ability to prevent the loss of control without requiring intensive offline computation besides superiority over the conventional state limiters in terms of maneuverability enhancement. The study’s conclusions demonstrate the significant promise of the proposed method, as it enables the aircraft to maintain stable behavior even subjected to harsh, excessive, and abrupt maneuver requests.
@@
The study is organized as follows: the modeling, employing over-actuated aerodynamic modeling and nonlinear equations of motion for the baseline over-actuated F-16, is briefly presented in Section 2. The flight control architecture is then separated into two distinct but related headlines: the design of the control augmentation system using incremental nonlinear dynamic inversion is given in Section 3.1, and the design of an optimization-based control allocation to handle the over-actuation is presented in Section 3.2. Afterwards, in Section 4, the attainable moment set and its advanced variation are presented, i.e. the incremental attainable moment set. At the core of the study, the architecture for the loss of control prevention, merging Lyapunov’s direct method and the inverse dynamics of the aircraft, is discussed in detail in Section 5. Furthermore, the required subsystems of the architecture are also presented in the same section. The proposed method’s superiority is demonstrated under distinct circumstances and aggressive maneuvers in Section 6. Additionally, a comparison between conventional state limiters and the proposed method is provided. Finally, the concluding remarks and presumable future works are shared in Section 7.
2 Nonlinear Flight Dynamics Model
@@
The baseline aircraft is the F-16 with an over-actuated control suite, meaning that it has five independent control surfaces: two ailerons (left and right), two horizontal tails (left and right), and one rudder. Hereby, the aerodynamic modeling must be adjusted in accordance with this control suite architecture by considering the contributions of each control surface separately. Nevertheless, the over-actuated aerodynamic modeling is adopted from [26, 27]. Each coefficient is expressed as a polynomial function involving the relevant states and control surface deflections. Also, the rigid body 6-degrees-of-freedom nonlinear equations of motion and necessary mass, inertia, actuator, and physical properties are given in [28]. As the cornerstone of this study, the nonlinear flight dynamics model is constructed based on the preceding studies in a high-fidelity manner.
@@ 3 Flight Control Architecture
In this section, the design of a control augmentation system using the incremental nonlinear dynamic inversion (INDI) with the fundamental time-scale separation principle is given. Additionally, an optimization-based control allocation design to deal with the over-actuation is discussed.
@@ 3.1 Control Augmentation System Design
In this study, the designed control augmentation system has three pilot inputs; angle of attack ($\alpha_{cmd}$), sideslip ($\beta_{cmd}$), and body-axis roll rate ($p_{cmd}$). The fast dynamics and slow dynamics are decomposed in accordance with the time-scale separation. The angle of attack and sideslip are the part of the slow dynamics, whereas the body angular rates are the part of fast dynamics. Therefore, the angle of attack and sideslip are the elements of the outer-loop, and the output of the outer-loop is used as the input of the inner-loop. As a consequence, the dynamics of $\alpha$ and $\beta$ can be such decomposed as the expression given in Eq. (1).
@@
where $\rho,\bar{c},b,S$ are the density, mean aerodynamic chord length, wing span, and wing area respectively. Also, $C_{L_{q}},C_{D_{q}},C_{Y_{r}}$ are the lift and drag coefficient derivatives with respect to the pitch rate, and side force coefficient derivative with respect to the yaw rate, respectively. Moreover, $[\dot{\alpha}\hskip 4.26773pt\dot{\beta}]^{T}$ is the virtual input and can be obtained using a simple proportional controller as given in Eq. (4).
@@
where $\omega_{\alpha}$ and $\omega_{\beta}$ are the proportional gains for $\alpha$ and $\beta$ channels, respectively. Eventually, the pilot inputs of $\alpha_{cmd}$ and $\beta_{cmd}$ are derived in the form of $q$ and $r$ as the inputs of the inner-loop.
@@
where $\omega_{p},\omega_{q},\omega_{r}$ are the proportional gains for $p,q$ and $r$ channels, respectively. Note that $q_{cmd}$ and $r_{cmd}$ are derived from the commanded angle of attack and sideslip in the outer loop, whereas $p_{cmd}$ is the direct pilot input. The proportional gains corresponding to each control channel in the proportional controllers are given in Table 1.

### v2

@@
Ege Ç. Altunkaya  [^1] Research Assistant, Aviation Institute, Aerospace Research Center, Istanbul Technical University, Istanbul, Türkiye, 34469; altunkaya16@itu.edu.tr and Akın Çatak  [^2] Research Assistant, Aviation Institute, Aerospace Research Center, Istanbul Technical University, Istanbul, Türkiye, 34469. Aerospace Research Center, Istanbul Technical University, Istanbul, Türkiye, 34469 Emre Koyuncu  [^3] Assoc. Prof., Aerospace Research Center, Istanbul Technical University, Istanbul, Türkiye, 34469. Aerospace Research Center, Istanbul Technical University, Istanbul, Türkiye, 34469 İbrahim Özkol  [^4] Prof., Aerospace Research Center, Istanbul Technical University, Istanbul, Türkiye, 34469. Aerospace Research Center, Istanbul Technical University, Istanbul, Türkiye, 34469
@@ 1 Introduction
The loss of control, often resulting from compromised stability and/or controllability, jeopardizes flight safety and is a leading factor in fatal crashes [1]. Numerous factors contribute to this loss of control, including adverse onboard conditions, such as system malfunctions or pilot error, and unexpected vehicle upsets caused by environmental disturbances or mechanical failures [2, 3]. The diverse array of reasons and their interactions inherently elevate the risk of control loss, as extensively outlined in terms of fatal crash statistics linked to the control loss in [2] and [3]. Moreover, a detailed analysis of the causes behind the control loss reveals that the pilot-induced loss of control, a significant contributing factor, involves abrupt and aggressive maneuvers that can trigger a destabilized flight [3]. Despite the desirability of performing aggressive maneuvers without compromising safety, the ability to do so remains constrained by the aircraft’s aerodynamic characteristics, controllability, and flight dynamics. Additionally, according to [4], reasons for loss of control can be categorized under five major headings: (1) stall, (2) sideslip-induced rolls, (3) rolls from other causes—such as wake vortices, structural failures, and lateral pilot-induced oscillations—, (4) pilot-induced oscillations, and (5) yaw departures—often resulting from wake encounters or improper rudder input. However, the loss of control is a broad concept that cannot be fully captured by simple analyses; nevertheless, a basic definition can be stated as follows: it is "a significant deviation of the aircraft from the intended flight path or operational envelope," as defined by the Commercial Aviation Safety Team (CAST) [5]. Moreover, CAST has proposed quantitative metrics to measure loss of control in [6]. Thereby designing an aircraft with a nonlinear flight control law is not always sufficient and does not imply that the aircraft can perform any desired of maneuver. There are numerous restrictions that hinder an agile aircraft from conducting agile maneuvers, such as high angle of attack kinematics, inertia-coupling, and flight and/or dynamic envelope violation. The factors described above may trigger an upset condition such as loss of control, departure, spin, stall, or even compromise the aircraft structural integrity. In summary, loss of control is a flight dynamics stability and control issue where the aircraft response deviates from controllable predictable conditions, involving non-linear effects like nonlinear kinematic and inertial coupling, which cause large amplitude responses to small changes in state variables or result in bounded oscillatory or divergent dynamics. As a consequence, it often leads to significant angular rates and displacements, making it challenging for the pilot to maintain control of the aircraft [7]. Therefore, to prevent any upset condition, one has to delve into presumable reasons for these catastrophic consequences.
@@
Since loss of control still persists as a critical issue, numerous studies inherently exist to address it from various perspectives. Hereby, ensuring safe flight has been extensively studied through flight envelope protection [8, 9, 10, 11, 12, 13, 14, 15, 16], prevention of loss of control [17, 11, 18], and implementation of command/reference governor methods [19, 20]. Lombaerts et al. [11] detail the design process of the adaptive algorithm and its integration into a piloted simulator for evaluation. The paper discusses the methodology employed in simulator testing, including the scenarios used to assess the algorithm’s performance. The boundaries of the flight envelope are determined adaptively, and the flight envelope-aware flight control system serves to prevent loss of control. Finally, they implement the proposed architecture into a simulator environment of a commercial airline aircraft, for further examination. Yu et al. [15] present an approach for designing safety control systems for aircraft that incorporate flight envelope protection and reference command generation. The approach is based on a nonlinear control framework that combines feedback control with feedforward reference commands to achieve robust and safe flight, accommodating actuator faults and ensuring operation within the predetermined safe flight envelope. The results show that the approach is effective at maintaining stable flight and protecting the aircraft from operating outside of its safe flight envelope. Despite the existence of adaptive flight envelope determination methods, there also exist other online safe flight envelope estimation studies [21, 22] besides the ANN-based flight envelope estimation approach [23].
@@
Nevertheless, various studies strongly emphasize the distinction between the flight envelope and dynamic envelope of an aircraft [24, 1]. In other words, the dynamic envelope defines the region within the aircraft’s state space where safe controllability is assured, preventing the occurrence of events leading to loss of control. Consequently, the notion of ensuring flight conditions solely within the traditional flight envelope for the prevention of the loss of control becomes a matter of debate. Particularly, Zheng et al. [24] study the determination of the dynamic envelope of the baseline aircraft, NASA’s Generic Transport Model. They generate the dynamic envelope using differential manifold theory and claim that this method is able to capture the coupling effects more accurately. Moreover, a correlation between the dynamic envelope and control coefficients of PI controller or pilot command of pitch angle, $\theta$ is established. Finally, based on the dynamic envelope enlargement strategy, an envelope protection architecture is proposed and assessed.
@@
On the other hand, various studies aim to estimate the maneuvering envelope through reachability analysis, such as [25, 26, 17, 22]. However, these approaches are not suitable for the real-time adaptation due to the high computational cost associated with numerically solving the Hamilton-Jacobi-Bellman equations. Alongside, the authors in [27] propose a real-time implementable online flight envelope detection mechanism that merges nonlinear reachability analysis with the classification capabilities of artificial neural networks. This online detection method is further utilized for accommodating structural failures in an aircraft, thereby preventing a potential loss of control, where the failure-induced loss of control is also a critical aspect and extensively addressed in [28]. Furthermore, in [1, 29], loss of control is thoroughly investigated in the sense of nonlinear dynamics, and bifurcation analysis is put forth. However, of course, bifurcation analyses can only be done in the design phase of the aircraft.
@@
Last but not least, maneuver limiting is also addressed in various studies using the command and reference governor approaches. Simon et al. [19] propose a method for limiting the angle of attack and load factor using the command and reference governors. The implementation has been carried out in the simulation environment used for the Jas39 Gripen fighter aircraft. In this study, both the angle of attack and load factor commands are saturated, leading to a significant reduction in overshoot in the angle of attack and load factor responses at their limit values by adjusting the command governor gain $\beta$, constrained by the corresponding flight envelope. As a consequence, since the predetermined safe flight envelope is protected, a loss of control is not observed during the desired maneuvers. A similar effort has been described in [20] for the flight envelope protection during high-angle-of-attack maneuvers. Albeit the proposed methods are successful in achieving their primary objectives, i.e., flight envelope protection, a priori flight envelope information is a must; otherwise, command and reference governors cannot determine the boundaries of the flight envelope in which the aircraft can be controlled without entering an upset condition.
@@
From an industrial perspective, the primary goal is "care-free" maneuvering without compromising agility and safety, as emphasized in [29]. However, aircraft certification remains a constant requirement in the industry, making conventional methods the preferred choice for safety. Traditional state limiters, such as angle of attack, load factor, bank angle, pitch angle, and roll rate limiters, are still in use. These limiters are scheduled based on Mach number, dynamic pressure, angle of attack, sideslip, and/or roll rate. Nevertheless, designing these state limiters for safe flight requires a labor-intensive effort. Extensive aerodynamic analyses are necessary to investigate the stability characteristics of the designed aircraft. For this purpose, analysis is typically conducted of several aircraft dynamics parameters, such as Lateral Control Departure Parameter (LCDP) and $C_{{n_{\beta}}_{dyn}}$. This analysis often involves varying the relevant states that define the limited envelope and using criteria to predict the aircraft’s tendency to departure based on the values of both the LCDP and $C_{{n_{\beta}}_{dyn}}$ parameters. Careful analysis of these parameters is of significant importance to determine the departure angle of attack value, and their variations with changing flight circumstances must be considered for scheduling. Importantly, these evaluations are not limited to the early design phases but must be extended and verified through the flight tests. Consequently, these aspects contribute to the labor and cost excessively.
@@
According to the inferences drawn from the previous research, as elaborated previously, the issue of the loss of control has been discussed within the context of the flight envelope protection. Additionally, previous studies often use pre-established flight envelope databases. As a result, the aircraft can be controlled with prior knowledge of critical values to avoid, based on the known limits of the flight envelope. However, confining the concept of the loss of control solely within the flight envelope protection does not seem inclusive, and the consideration of extreme and coupling maneuvers has been lacking in most previous research.
1.1 Objective, Methodology, and Contributions
This study aims to develop an online maneuver limiting strategy rather than a flight envelope protection system, with the goal of performing the most extreme maneuvers possible—beyond the limits imposed by conventional state limiters. Therefore, the goal is to perform maneuvers as extreme as possible by limiting only the pilot commands, rather than structural or attitude features such as load factors, airspeed, and angular orientations. To achieve this, a two-step strategy is proposed: (1) online loss of control detection using an incremental attainable moment set, and (2) online loss of control prevention using pilot command limiting with a Lyapunov-based command limiter. Based on these evaluations, the followings are the contributions of this study:
  • This study does not use pre-existing a priori information for defining avoidance limits. Instead, it focuses on ensuring maneuver stability by intervening and limiting only the pilot commands for angle of attack ($\alpha$), roll rate ($p$), pitch rate ($q$), and yaw rate ($r$) in-flight. This process utilizes a combination of Lyapunov’s direct method and inverse dynamics.
  • The concept of incremental attainable moment set and its derivation are introduced for the first time in the investigation of aircraft controllability.
  • The application of these moment sets for assessing controllability in real-time is discussed.
  • Extreme maneuvers are examined to explore the aircraft’s limits, taking into account coupling and nonlinear effects. This approach contrasts with previous studies, and the effectiveness of the proposed method is demonstrated through these rigorous maneuvers.
@@ 1.1 Objective, Methodology, and Contributions
  • A comparison between conventional state limiters and the proposed method reveals significantly larger volumes of stable maneuverability. This result highlights the improvement in the aircraft’s agility, showing enhanced maneuverability without compromising safety.
@@
Consequently, the importance of the proposed method lies in its ability to prevent loss of control without requiring intensive offline computation and in its higher effectiveness in preserving aircraft maneuverability, with respect to the conventional state limiters. The conclusions of the study demonstrate the significant promise of the proposed method, as it enables the pilot to retain control of the aircraft in the execution of large amplitude and abrupt maneuvers.
@@
The study is structured as follows: Section 2 establishes the necessary background, covering 2.1 for the modeling of the over-actuated F-16 as a baseline, and 2.2 the introduction of the incremental attainable moment set for the loss of control detection. Afterwards, Section 3.1 covers the design of the control augmentation system using incremental nonlinear dynamic inversion, while Section 3.2 discusses the optimization-based control allocation for over-actuation. The core architecture for loss of control prevention, combining Lyapunov’s direct method and aircraft inverse dynamics, is detailed in Section 4, along with its subsystems. The effectiveness of the proposed method is demonstrated through various scenarios and aggressive maneuvers, comparing it to conventional state limiters in Section 5. Finally, Section 6 provides concluding remarks and potential future work.
2 Preliminaries
In this section, the necessary background is established: (1) the nonlinear flight dynamics model as a baseline, and (2) the incremental attainable moment set concept, which will be leveraged for loss of control detection, as discussed in the proceeding sections.
2.1 Nonlinear Flight Dynamics Model
@@
The reference aircraft for the simulations is the F-16 fighter jet, equipped with an over-actuated control suite, i.e. five independent control surfaces—two ailerons (left and right), two horizontal tails (left and right), and one rudder—each capable of independent deflection directions, magnitudes, and rates based on the desired control action. Hereby, the aerodynamic modeling must be adjusted in accordance with this control suite architecture by considering the contributions of each control surface separately. Nevertheless, the over-actuated aerodynamic modeling is adopted from [30, 31]. Each aerodynamic coefficient is expressed as a polynomial function involving the relevant states and control surface deflections. Also, the rigid body 6-degrees-of-freedom nonlinear equations of motion and necessary mass, inertia, actuator, and physical properties are given in [32]. As the cornerstone of this study, the nonlinear flight dynamics model is constructed based on the preceding studies in a high-fidelity manner.
2.2 Incremental Attainable Moment Set
The attainable moment set (AMS) of an aircraft is a geometrical representation (topologically defines a convex-hull) of the aerodynamic moments that can be generated by the aircraft’s control effectors at various flight conditions. Furthermore, an attainable moment set of an aircraft helps understanding the instantaneous control authority of the aircraft, i.e. controllability of the aircraft [33, 34]. This direct correlation can be established under the following conditions: omitting (1) inertial couplings, (2) dynamic pressure, and (3) external factors such as atmospheric disturbances. Since these factors are not functions of the aerodynamic moment coefficients, the investigation of controllability can be distilled directly to control authority. Due to this benefit, the AMS will be used in this study to detect the possibility of loss of control. However, constructing the attainable moment set based solely on actuator position limits results in the globally achievable control moment coefficient under the specified circumstances. Instead, an alternative path is proposed to give a better insight about the instantaneous controllability during an agile maneuver, i.e. incremental attainable moment set.
The concept and utilization of the incremental attainable moment set (IAMS) is the same as the attainable moment set with a distinction: the boundary formulation. In the proposed method, both the actuator deflection limits and the actuator rate limits are considered to determine whether the aircraft can provide the necessary control moment at each time step or not. Therefore, with a slight notion shift, the attainable moment set is reformulated to construct the incremental attainable moment set $\Delta\mathbb{A_{M}}$ as given in Eq. (1).
@@ 3 Flight Control Architecture
This section contains the description of the design of a control augmentation system using the incremental nonlinear dynamic inversion (INDI) with the fundamental time-scale separation principle, which is a commonly preferred method [35, 36, 37]. Additionally, an optimization-based control allocation design to deal with the over-actuation is discussed.
@@ 3.1 Control Augmentation System Design
In this study, the designed control augmentation system has three pilot commands; angle of attack ($\alpha_{cmd}$), sideslip ($\beta_{cmd}$), and body-axis roll rate ($p_{cmd}$). Such a control output scheme is regarded as more suitable for managing extreme maneuvers, as aerodynamic angles provide more information about the aircraft’s aerodynamic characteristics than angular rates do. However, it is important to note that angle of attack and sideslip sensors are relatively low-rate and sensitive to atmospheric disturbances or aerodynamic interferences, which may lead to a deterioration in the performance of INDI control.
The fast dynamics and slow dynamics are decomposed in accordance with the time-scale separation. The angle of attack and sideslip are the part of the slow dynamics, whereas the body angular rates are the part of fast dynamics. Therefore, the angle of attack and sideslip are the elements of the outer-loop, and the output of the outer-loop is used as the input of the inner-loop. As a consequence, the dynamics of $\alpha$ and $\beta$ can be such decomposed as the expression given in Eq. (3).
@@
where $\rho,\bar{c},b,S$ are the density, the mean aerodynamic chord length, the wing span, and the wing area respectively. Also, $C_{L_{q}},C_{D_{q}},C_{Y_{r}}$ are the lift and drag coefficient derivatives with respect to the non-dimensional pitch rate, and side force coefficient derivative with respect to the non-dimensional yaw rate, respectively. Moreover, $[\dot{\alpha}\hskip 4.26773pt\dot{\beta}]^{T}$ is the virtual input and can be obtained using a simple proportional controller as given in Eq. (6).
@@
where $\omega_{\alpha}$ and $\omega_{\beta}$ are the proportional gains for $\alpha$ and $\beta$ channels, respectively. Eventually, the pilot commands for angle of attack ($\alpha_{cmd}$) and sideslip angle ($\beta_{cmd}$) are converted into desired angular rates, specifically roll rate ($p$), pitch rate ($q$), and yaw rate ($r$), which serve as inputs for the inner-loop control system, which will be illustrated in Fig 1.
@@
where $\omega_{p},\omega_{q},\omega_{r}$ are the proportional gains for $p,q$ and $r$ channels, respectively. Note that $q_{cmd}$ and $r_{cmd}$ are derived from the commanded angle of attack and sideslip in the outer loop, whereas $p_{cmd}$ is the direct pilot commands. The proportional gains corresponding to each control channel in the proportional controllers are given in Table 1.